# The Gorini-Kossakowski-Sudarshan-Lindblad problem and the geometry of CP maps

### v4

@@ 10.3.2 Some properties of $\bm{\theta}$
(c). $d{\bm{\theta}}(A)\cdot B=B\,\square\,A^{\dagger}+A\,\square\,B^{\dagger}$ is a bounded $\mathbb{R}$}-linear operator. So, $\bm{\theta}$} is not only norm-continuous, but norm-differentiable. By part (b), $\mathscr{P}_{\alpha}({\bm{\theta}}(A))={\bm{\theta}}(\mathscr{P}_{\alpha}(A))$. Thus, by the just-proven norm continuity, dd/$d$ continuity follows. Finally, if $\rho=\sum_{i=1}^{\infty}{\bm{\pi}}(h_{i})$, then for $B\in\overline{\mathrm{ball}}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$, $B\rho B^{\dagger}\approx A\rho A^{\dagger}$ as soon as $Bh_{i}\approx Ah_{i}$ for $i=1,\ldots,N$ for some $N$. This shows SOT/SOT continuity on bounded sets. ∎
@@ 10.4 Extreme CP map = $\bm{\theta}$-map
If $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ is not an extreme CP map, then for sufficiently large $\alpha$, $\mathscr{P}_{\alpha}\Lambda$ is not extreme.
@@
Suppose $\Lambda=\Gamma+\Gamma^{\prime}$ is a non-trivial CP decomposition. Then, $\mathscr{P}_{\alpha}\Lambda=\mathscr{P}_{\alpha}\Gamma+\mathscr{P}_{\alpha}\Gamma^{\prime}$; the terms on the right-hand side tend to $\Gamma$ and $\Gamma^{\prime}$, respectively, in SOT. Now, suppose for a contradiction that along some subsequence, $\mathscr{P}_{\alpha}\Gamma=c_{\alpha}\mathscr{P}_{\alpha}\Gamma^{\prime}$ with $0<c_{\alpha}<\infty$. Then by SOT convergence, it is not possible that the $c_{\alpha}$ are unbounded. Therefore, there is a subsequence along which $c_{\alpha}$ converges to $c\geq 0$. But then, $\Gamma=c\Gamma^{\prime}$. ∎
@@
Given $0\prec\Lambda$, there is $h$ such that $\Lambda{\bm{\pi}}(h)$ has a nontrivial orthogonal decomposition ${\bm{\pi}}(k)+\cdots$. Take some filtration with $\mathcal{H}_{1}=\operatorname{\mathrm{span}}\{h,k\}$. Then, $(\mathscr{P}_{\alpha}\Lambda){\bm{\pi}}(h)=\mathscr{P}_{\alpha}(\Lambda{\bm{\pi}}(h))={\bm{\pi}}(k)+\cdots$. By Lemma 10.12 below, Kraus decompositions $\mathscr{P}_{\alpha}\Lambda=\sum_{i}{\bm{\theta}}(A^{\alpha}_{i})$ of the finite-dimensional projections of $\Lambda$ may be chosen so that $\left\langle{k}\middle|{A_{1}^{\alpha}h}\right\rangle=\|k\|$. Since $\Lambda\succsim{\bm{\theta}}(A_{1}^{\alpha})$, the sequence $(A_{1}^{\alpha})$ is bounded and therefore has a $d$ cluster point, $A$, which is nonzero because $\left\langle{k}\middle|{Ah}\right\rangle=\|k\|$. Summing up, $\Lambda\succsim{\bm{\theta}}(A)\succ 0$. ∎
Lemma 10.12.
Let $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ where $\mathcal{H}$, $\mathcal{K}$ are finite-dimensional, and let $\Lambda{\bm{\pi}}(h)$ have orthogonal decomposition $\Lambda{\bm{\pi}}(h)={\bm{\pi}}(k)+\cdots$. Then there exists a choice of $A_{i}$ in the Kraus decomposition $\Lambda=\sum_{i}{\bm{\theta}}(A_{i})$ such that
$\left\langle{k}\middle|{A_{i}h}\right\rangle=\begin{cases}\|k\|,&i=1\\ 0,&\text{otherwise}.\end{cases}$ (10.3)
Proof.
For convenience, scale $h$ so that $\|k\|=1$. Let $\Lambda=\sum_{j}{\bm{\theta}}(B_{j})$ be any Kraus decomposition. Then
@@ 10.5 Order closedness of $\mathsf{CP}(\mathcal{H},\mathcal{K})$
A $\leq$-directed set $D$ in $\mathcal{B}^{1}(\mathcal{H})$ which is bounded above has a supremum and converges to it in norm.
@@
The overall strategy of the proof is much like that of Lemma 10.13. Given $\rho\colon\mathcal{B}^{1}(\mathcal{H})$, $D(\rho){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\{\Gamma\rho\,:\,\Gamma\in D\}$ is a directed set in $\mathcal{B}^{1}(\mathcal{H})$, bounded above, exactly the subject of Lemma 10.13. Define a function $f(\rho){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\sup(D(\rho))$; $D(\rho)$ converges to $\sup(D(\rho))$ in norm, so $D$ converges to $f$ in SOT. It only remains to used directedness to show that $f$ is linear: Given $\rho,\rho^{\prime}$, there is $\Lambda\in D$ such that $f(\rho+c\sigma)\approx\Lambda(\rho+c\sigma)=\Lambda\rho+c\Lambda\sigma\approx f(\rho)+cf(\sigma)$. ∎
@@ 10.6 Every CP map has Kraus decompositions
Prop. 10.15 shows that $\Lambda_{\infty}=\lim\Lambda_{n}$ and $K_{\infty}=\lim K_{n}$ exist in $\mathsf{CP}(\mathcal{H},\mathcal{K})$. Suppose, for a contradiction, that $0\prec\Lambda_{\infty}$. Then, by Lemma 10.11, as usual, there is some $0\prec{\bm{\theta}}(A)\precsim\Lambda_{\infty}$, and $f({\bm{\theta}}(A))>0$. But $A$ belongs to every $S_{n}$, and this is impossible by (10.6). Therefore, $K_{\infty}=\Lambda$ and the $A_{n}$ are the elements of a Kraus decomposition.
@@ 11 GKSL problem in separable setting
11.1 Dimension-independent bound on the minimal Lindblad parametrizer
@@ 11.3 Lindblad parametrizations and parametrizers exist
Because $\mathscr{P}_{\alpha}$ is a contraction, if $\Lambda$ is forward tangent to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathcal{H}}$, then $\mathscr{P}_{\alpha}\Lambda$ is forward tangent to $\mathsf{CP}(\mathcal{H}_{\alpha})$ at $\operatorname{\mathbbm{1}}_{\mathcal{H}_{\alpha}}$. Therefore, as an element of $HP(\mathcal{H}_{\alpha})\times\mathrm{SA}(\mathcal{H}_{\alpha})\times\mathrm{SA}(\mathcal{H}_{\alpha})$, $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ has CP first component. But it is still CP regarded as an element of $\mathsf{HP}(\mathcal{H})$. ∎
@@ 11.3.2 The sequence $(\Delta^{\scriptstyle{min}}_{\alpha})$ is an approximate right-inverse of $L$
Now, since (Lemma 11.1) $L$ and $\Delta^{\scriptstyle{min}}_{\alpha}$ are bounded uniformly in $\alpha$, it suffices to show that $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)\cdot A\rightarrow\Lambda A$ for $A$ in the set $\cup_{m}\mathcal{B}^{1}(\mathcal{H}_{m})$ dense in $\mathcal{B}^{1}(\mathcal{H})$. Thus, take AA in $\mathcal{B}^{1}(\mathcal{H}_{\beta})$. As soon as $\beta\leq\alpha$, $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)A=(\mathscr{P}_{\alpha}\Lambda)A$, by (11.10). But, $\mathscr{P}_{\alpha}\Lambda\xrightarrow{SOT}\Lambda$, and the conclusion follows immediately. ∎
11.3.3 Bringing in the $d$-metrics
Eventually, we must work with the $d$-metrics because that is the only topology we have with respect to which every bounded sequence has cluster points.
Lemma 11.4.
$L\colon\mathsf{Lind\,{\mathcal{H}}}\rightarrow\mathsf{HP}(\mathcal{H})$ is dd/$d$ continuous.
@@ 11.3.3 Bringing in the $d$-metrics
Let $Z_{n}\xrightarrow{d}Z$ in $\mathsf{Lind\,{\mathcal{H}}}$. By definition (of $d$-convergence), this means that $\mathscr{P}_{\alpha}(Z_{n})\to\mathscr{P}_{\alpha}(Z)$ for each $\alpha$. Then, applying $L$ and composing with $\mathscr{P}_{\alpha}$, $L(\mathscr{P}_{\alpha}(Z_{n}))\circ\mathscr{P}_{\alpha}\to L(\mathscr{P}_{\alpha}(Z))\circ\mathscr{P}_{\alpha}$. Lemma 11.8c shows that this is the same as $\mathscr{P}_{\alpha}(LZ_{n})\to\mathscr{P}_{\alpha}(LZ)$, which, again by definition, is just $LZ_{n}\xrightarrow{d}LZ$. Therefore $Z$ is a Lindblad parametrization.
It remains to show that the first component of $Z$ is CP if $\Lambda\in\mathsf{cp}^{+}(\mathcal{H})$. This follows immediately from $d$-closure of $\mathsf{CP}(\mathcal{H})$ (Prop. 10.5) and the fact that the first component of each $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ is CP. ∎
@@ 11.3.4 Individual Lindblad parametrizations
$(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)$ is bounded by Lemma 11.1, hence has $d$-cluster-points (see §9). Let $Z$ be one such. By Lemma 11.4, $LZ$ is a $d$-cluster point of the sequence $(L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda))$. But Lemma 11.3 says that only $\Lambda$ can be such a cluster point.
@@ 11.3.5 Full Lindblad parametrizer
In $\mathsf{Lind\,{\mathcal{H}}}$ equipped with the $d$-metric, $\left\{\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n}\,\colon\,\alpha\in{\mathbb{N}}\right\}$ is a precompact set. The product space $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$ can be turned into a metric space just as in §9, such that products of compact sets are compact. Thus, the sequence $\alpha\mapsto\prod_{n}\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n}$ is precompact in $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$. $L^{\mathbb{N}}$ carries any cluster point of it into a cluster point of the sequence $\alpha\mapsto\prod_{n}L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n})$ in $\mathsf{HP}(\mathcal{H})^{\mathbb{N}}$, by Lemma 11.4.
@@
To that end, suppose $(\Lambda_{n})$ is a sequence in $\mathsf{cp}^{+}(\mathcal{H})$, $d$-convergent to, say, $\Lambda$. In addition, assume that this sequence is bounded in norm. Later, we will see how to remove this restriction. By Prop. 11.6, or even Lemma 11.5, there is a bounded sequence $Z_{n}$ of Lindblad parametrizations in $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. This latter sequence has $d$-cluster points; let $Z$ be one such. Necessarily, it too is in $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. By Lemma 11.4, $LZ$ is a $d$-cluster point of $(\Lambda_{n})$. Since $\Lambda_{n}\xrightarrow{d}\Lambda$, it follows that $LZ=\Lambda$.

### v5

@@ 10.3.2 Some properties of $\bm{\theta}$
(c). $d{\bm{\theta}}(A)\cdot B=B\,\square\,A^{\dagger}+A\,\square\,B^{\dagger}$ is a bounded $\mathbb{R}$}-linear operator. So, $\bm{\theta}$} is not only norm-continuous, but norm-differentiable. Then, $\bm{\theta}$} is dd/$d$-continuous by part (b) and Lemma 9.3. Finally, if $\rho=\sum_{i=1}^{\infty}{\bm{\pi}}(h_{i})$, then for $B\in\overline{\mathrm{ball}}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$, $B\rho B^{\dagger}\approx A\rho A^{\dagger}$ as soon as $Bh_{i}\approx Ah_{i}$ for $i=1,\ldots,N$ for some $N$. This shows SOT/SOT continuity on bounded sets. ∎
@@ 10.4 Extreme CP map = $\bm{\theta}$-map
If $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ is not an extreme CP map, then for sufficiently large $\alpha$, $\mathscr{P}_{\alpha}\Lambda$ is not extreme. Equivalently, if $\Lambda$ is extreme, then so is every projection $\mathscr{P}_{\alpha}\Lambda$.
@@
Suppose $\Lambda=\Gamma+\Gamma^{\prime}$ is a non-trivial CP decomposition. Then, $\mathscr{P}_{\alpha}\Lambda=\mathscr{P}_{\alpha}\Gamma+\mathscr{P}_{\alpha}\Gamma^{\prime}$; the terms on the right-hand side tend in SOT to $\Gamma$ and $\Gamma^{\prime}$, respectively, by Example 8.5. Now, suppose for a contradiction that along some subsequence, $\mathscr{P}_{\alpha}\Gamma=c_{\alpha}\mathscr{P}_{\alpha}\Gamma^{\prime}$ with $0<c_{\alpha}<\infty$. Then by SOT convergence, it is not possible that the $c_{\alpha}$ are unbounded. Therefore, there is a subsequence along which $c_{\alpha}$ converges to $c\geq 0$. But then, $\Gamma=c\Gamma^{\prime}$. ∎
Corollary 10.10.
The set of extreme CP maps in $\mathsf{CP}(\mathcal{H},\mathcal{K})$ is $d$-closed.
Proof.
Suppose $(\Lambda_{n})$ is a sequence of extreme CP maps and $\Lambda_{n}\xrightarrow{d}\Lambda$, that is, $\mathscr{P}_{\alpha}\Lambda_{n}\xrightarrow{\mathscr{P}}_{\alpha}\Lambda$ for each $\alpha$. By the preceding Lemma, each $\mathscr{P}_{\alpha}\Lambda_{n}$ is extreme, hence so is $\mathscr{P}_{\alpha}\Lambda$ by the finite dimensional theory (§4.2.2). Therefore, again by the preceding Lemma, $\Lambda$ is extreme. ∎
@@
In proving this, the following finite-dimensional lemma will be instrumental, for which one should recall the discussion of §4.2.1.
Lemma 10.13.
Let $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ where $\mathcal{H}$, $\mathcal{K}$ are finite-dimensional, and let $\Lambda({\bm{\pi}}(h))$ have orthogonal decomposition $\Lambda{\bm{\pi}}(h)={\bm{\pi}}(k)+\cdots$. Then there exists a choice of $A_{i}$ in the Kraus decomposition $\Lambda=\sum_{i}{\bm{\theta}}(A_{i})$ such that
$\left\langle{k}\middle|{A_{i}h}\right\rangle=\begin{cases}\|k\|\|h\|,&i=1\\ 0,&\text{otherwise}.\end{cases}$ (10.3)
@@
For convenience, scale $h$ so that $\left\langle{{\bm{\pi}}(k)}\middle|{\Lambda{\bm{\pi}}(h)}\right\rangle=1$. Let $\Lambda=\sum_{j}{\bm{\theta}}(B_{j})$ be any Kraus decomposition. Then
@@
Proof of Lemma 10.12.
Given $0\prec\Lambda$, there is $h$ such that $\Lambda{\bm{\pi}}(h)$ has a nontrivial orthogonal decomposition ${\bm{\pi}}(k)+\cdots$. Take some filtration with $\mathcal{H}_{1}=\operatorname{\mathrm{span}}\{h,k\}$. Then, $(\mathscr{P}_{\alpha}\Lambda){\bm{\pi}}(h)=\mathscr{P}_{\alpha}(\Lambda{\bm{\pi}}(h))={\bm{\pi}}(k)+\cdots$. By Lemma 10.13 below, Kraus decompositions $\mathscr{P}_{\alpha}\Lambda=\sum_{i}{\bm{\theta}}(A^{\alpha}_{i})$ of the finite-dimensional projections of $\Lambda$ may be chosen so that $\left\langle{k}\middle|{A_{1}^{\alpha}h}\right\rangle=\|k\|$. Since $\Lambda\succsim{\bm{\theta}}(A_{1}^{\alpha})$, the sequence $(A_{1}^{\alpha})$ is bounded and therefore has a $d$ cluster point, $A$, which is nonzero because $\left\langle{k}\middle|{Ah}\right\rangle=\|k\|$. Since $\bm{\theta}$} is $d/d$ continuous (Lemma 10.8, ${\bm{\theta}}(A_{1}^{\alpha})\xrightarrow{d}{\bm{\theta}}(A)\succ 0$, and then by Prop. 10.6, $\Lambda\succsim{\bm{\theta}}(A)$. ∎
@@ 10.5 Order closedness of $\mathsf{CP}(\mathcal{H},\mathcal{K})$
A $\leq$-directed set $D$ in $\mathfrak{B}^{1}(\mathcal{H})$ which is bounded above has a supremum and converges to it in norm.
@@
The overall strategy of the proof is much like that of Lemma 10.14. Given $\rho\colon\mathfrak{B}^{1}(\mathcal{H})$, $D(\rho){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\{\Gamma\rho\,:\,\Gamma\in D\}$ is a directed set in $\mathfrak{B}^{1}(\mathcal{H})$, bounded above, exactly the subject of Lemma 10.14. Define a function $f(\rho){\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\sup(D(\rho))$; $D(\rho)$ converges to $\sup(D(\rho))$ in norm, so $D$ converges to $f$ in SOT. It only remains to used directedness to show that $f$ is linear: Given $\rho,\rho^{\prime}$, there is $\Lambda\in D$ such that $f(\rho+c\sigma)\approx\Lambda(\rho+c\sigma)=\Lambda\rho+c\Lambda\sigma\approx f(\rho)+cf(\sigma)$. ∎
@@ 10.6 Every CP map has Kraus decompositions
Prop. 10.16 shows that $\Lambda_{\infty}=\lim\Lambda_{n}$ and $K_{\infty}=\lim K_{n}$ exist in $\mathsf{CP}(\mathcal{H},\mathcal{K})$. Suppose, for a contradiction, that $0\prec\Lambda_{\infty}$. Then, by Lemma 10.12, as usual, there is some $0\prec{\bm{\theta}}(A)\precsim\Lambda_{\infty}$, and $f({\bm{\theta}}(A))>0$. But $A$ belongs to every $S_{n}$, and this is impossible by (10.6). Therefore, $K_{\infty}=\Lambda$ and the $A_{n}$ are the elements of a Kraus decomposition.
@@ 11 GKSL problem in separable setting
11.1 Dimension-independent bound on minimal Lindblad parametrizers
@@ 11.3 Lindblad parametrizations and parametrizers exist
Because $\mathscr{P}_{\alpha}$ is a contraction, if $\Lambda$ is forward tangent to $\mathsf{CP}(\mathcal{H})$ at $\operatorname{\mathbbm{1}}_{\mathcal{H}}$, then $\mathscr{P}_{\alpha}\Lambda$ is forward tangent to $\mathsf{CP}(\mathcal{H}_{\alpha})$ at $\operatorname{\mathbbm{1}}_{\mathcal{H}_{\alpha}}$. Therefore, as an element of $\mathsf{HP}(\mathcal{H}_{\alpha})\times\mathrm{SA}(\mathcal{H}_{\alpha})\times\mathrm{SA}(\mathcal{H}_{\alpha})$, $\Delta^{\scriptstyle{min}}_{\alpha}\Lambda$ has CP first component. But it is still CP regarded as an element of $\mathsf{HP}(\mathcal{H})$. ∎
@@ 11.3.2 The sequence $(\Delta^{\scriptstyle{min}}_{\alpha})$ is an approximate right-inverse of $L$
Now, since (Lemma 11.1) $L$ and $\Delta^{\scriptstyle{min}}_{\alpha}$ are bounded uniformly in $\alpha$, it suffices to show that $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)\cdot A\rightarrow\Lambda A$ for $A$ in the set $\cup_{m}\mathfrak{B}^{1}(\mathcal{H}_{m})$ dense in $\mathfrak{B}^{1}(\mathcal{H})$. Thus, take AA in $\mathfrak{B}^{1}(\mathcal{H}_{\beta})$. As soon as $\beta\leq\alpha$, $L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)A=(\mathscr{P}_{\alpha}\Lambda)A$, by (11.10). But, $\mathscr{P}_{\alpha}\Lambda\xrightarrow{SOT}\Lambda$, and the conclusion follows immediately. ∎
11.3.3 $L$ is $d/d$-continuous
Eventually, we must work with the $d$-metrics because that is the only topology we have with respect to which any bounded sequence has cluster points.
Proposition 11.4.
$L\colon\mathsf{Lind\,{\mathcal{H}}}\rightarrow\mathsf{HP}(\mathcal{H})$ is dd/$d$-continuous.
@@ 11.3.3 $L$ is $d/d$-continuous
Let $Z_{n}\xrightarrow{d}Z$ in $\mathsf{Lind\,{\mathcal{H}}}$. By definition (of $d$-convergence), this means that $\mathscr{P}_{\alpha}(Z_{n})\to\mathscr{P}_{\alpha}(Z)$ for each $\alpha$. Then, applying $L$ and composing with $\mathscr{P}_{\alpha}$, $L(\mathscr{P}_{\alpha}(Z_{n}))\circ\mathscr{P}_{\alpha}\to L(\mathscr{P}_{\alpha}(Z))\circ\mathscr{P}_{\alpha}$. This is the same as $\mathscr{P}_{\alpha}(LZ_{n})\to\mathscr{P}_{\alpha}(LZ)$ by (11.8c), which, again by definition, is just $LZ_{n}\xrightarrow{d}LZ$. ∎
@@ 11.3.4 Individual Lindblad parametrizations
$(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda)$ is bounded by Lemma 11.1, hence has $d$-cluster-points (see §9). Let $Z$ be one such. By Prop. 11.4, $LZ$ is a $d$-cluster point of the sequence $(L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda))$. But Lemma 11.3 says that only $\Lambda$ can be such a cluster point.
@@ 11.3.5 Full Lindblad parametrizer
In $\mathsf{Lind\,{\mathcal{H}}}$ equipped with the $d$-metric, $\left\{\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n}\,\colon\,\alpha\in{\mathbb{N}}\right\}$ is a precompact set. The product space $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$ can be turned into a metric space just as in §9, such that products of compact sets are compact. Thus, the sequence $\alpha\mapsto\prod_{n}\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n}$ is precompact in $(\mathsf{Lind\,{\mathcal{H}}})^{\mathbb{N}}$. $L^{\mathbb{N}}$ carries any cluster point of it into a cluster point of the sequence $\alpha\mapsto\prod_{n}L(\Delta^{\scriptstyle{min}}_{\alpha}\Lambda_{n})$ in $\mathsf{HP}(\mathcal{H})^{\mathbb{N}}$, by Prop. 11.4.
@@
To that end, suppose $(\Lambda_{n})$ is a sequence in $\mathsf{cp}^{+}(\mathcal{H})$, $d$-convergent to, say, $\Lambda$. In addition, assume that this sequence is bounded in norm. Later, we will see how to remove this restriction. By Prop. 11.6, or even Lemma 11.5, there is a bounded sequence $Z_{n}$ of Lindblad parametrizations in $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. This latter sequence has $d$-cluster points; let $Z$ be one such. Necessarily, it too is in $\mathsf{CP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})/{\mathbb{R}}$. By Prop. 11.4, $LZ$ is a $d$-cluster point of $(\Lambda_{n})$. Since $\Lambda_{n}\xrightarrow{d}\Lambda$, it follows that $LZ=\Lambda$.